In an experiment, positive and negative values are equally likely to occur. The probability of obtaining at most one neg
Think of it like this: p=0.5, n=5. P(at most 1 negative)=P0+P1= C5,0(0.5)^5+C5,1(0.5)^5=1/32+5/32=6/32=3/16. Full.
Ref: NCERT Mathematics Class XII — Probability